# A Data-Driven Approach for Topology Correction in Low Voltage Networks with DERs

### v1

@@ Abstract
This paper introduces a data-driven topology identification and correction approach for low-voltage distribution networks (LVDNs) combined with a time-based smart meter data selection strategy, aiming to correct outdated recordings and identify the missed recordings. The proposed approach solely relies on voltage magnitude measurements, releasing privacy concerns and measurement burdens. It enables the distribution system operators to identify switch states through supervised learning algorithms, as well as determine user-feeder connections and phase labels of customers by a modified Hierarchical Clustering algorithm. To address the similarity among smart meter (SM) data caused by distributed photovoltaic (PV) systems, a time-based SM data selection strategy is combined with the proposed correlation analysis. The feasibility and robustness of the proposed approach are validated using modified real-world LVDNs and multiple incomplete SM datasets collected from customers in the Netherlands. The results demonstrate that the time-based SM data selection strategy effectively mitigates their impact on phase identification, and the corrected topology not only improves network observability but also supports network operators in load balancing and PV consumption.
Index Terms:
Distribution networks, Topology observability, supervised learning, Unbalanced networks, Robustness
I Introduction
@@
Balanced distribution networks contributes to reducing voltage violations, line congestion, power loss, etc. Most existing phase balancing and reconfiguration problems are formulated as mixed-integer optimization problems that depend on network topologies [1]. However, these topologies are often unavailable or inaccurate for distribution system operators (DSOs) due to outdated or missing data, such as limited phase labels [2]. The increasing uncertainty of distributed energy resources (DERs), including household photovoltaic (PV) systems, heating pumps etc., impacts the topology reconfiguration frequency and challenges the identification of low-voltage distribution network (LVDN) topology [3, 4, 5]. Moreover, existing topology identification methods rely on complete smart meter (SM) datasets, the availability of which is often limited due to privacy concerns [6, 7]. Synthetic networks and benchmark models in[8, 9] are not practical to represent European LVDNs. Thus, practical topology correction methods for LVDNs are needed for active management and DERs penetration.
@@
The topology identification problem in distribution networks (DNs) could be briefly classified into three categories based on the main issues they are aiming to solve: (1) switch state identification, (2) connection lines and their length identification and (3) phase identification. The switch state identification was modeled as a classification problem in [10], solved by a trained deep neural network and providing quick estimations under the selected features. To recover the accurate connection lines between transformers and customers, a probabilistic graphical model-based method was developed in [11] by restricting the propagation of SM data error. A power band-based strategy is proposed for dataset selection, aiming to enhance the accuracy of phase identification and transformer-to-meter matching [12]. In [13], a hybrid data-driven approach integrating a partial correlation analysis strategy and a linear regression model is proposed to estimate topology from limited SM data. A similar work in [14] established a data-driven model based on the impedance matrix to correct topology and identify the impedance of connection lines from SM data.
@@
Based on an alternating direction method of multipliers, a robust topology identification method is proposed to estimate the line connection and their parameters using $\mu$-PMUs and SM datasets [15]. Based on incomplete and asynchronous SM data, an optimization-based data-driven approach is constructed to recover the approximate topologies in LVDNs with open datasets [16]. A two-step approach is introduced to identify the topology of LVDNs with multiple meter types, which provide different resolution datasets, mitigating the dependence on the SMs [17]. For networks with low observability, a distance matrix is introduced in [18] to correct the topology and is also effective in detecting latent nodes positioned between observed nodes along network paths (i.e., intermediate latent nodes rather than endpoints). A physics-informed neural network (PINN) is constructed to estimate the topology by representing the nodes without meters, whose parameters are adjusted by a reinforcement learning-based approach [19]. Although the above methods generate accurate topologies of DNs, the requirements of complete SM data and prior topology information (e.g., the topology candidates) make them infeasible in practice due to data privacy-related concerns. Moreover, optimizing the line resistance and reactance as independent variables may lead to unrealistic parameter estimates, as the properties are inherently interdependent and influenced by physical characteristics of the conductors and environments, such as temperature, humidity.
@@
Related to phase identification, current approaches rely on voltage magnitude data and demand profiles [20]. Since the correlation of time-series voltage magnitude data from the same phase is stronger than that from different phases, correlation analysis is normally used. Given voltage magnitude datasets, spectral clustering and $k$-means clustering approaches are adopted to identify the customer phase label [21]. The mixed integer linear programming model in [22] is constructed to identify the phase labels of single-phase customers by minimizing the gap between transformers’ measurement and the estimated data (e.g., power and voltage magnitude) in each phase. An optimization model-based approach was introduced in[1] to estimate customer phase labels in LVDNs with limited available recordings. Saliency analysis (SA) was integrated to extract the underlying features aiming to enhance the accuracy of phase label identification [23] SA and statistical tests are combined in[24] to guarantee the identification accuracy of the data pairs with weak correlations. Voltage magnitude fluctuations among customers connected to different phases may exhibit similarities due to variations induced by the medium-voltage (MV) network or similar load patterns. To mitigate this effect, a high-pass filter is applied during time-series data pre-processing in the phase identification approach[25]. Beyond hard classification (i.e., assigning customers a binary label of 0 or 1), soft classification is also considered for phase identification. The method in[26] assigns three probability values to customers. The highest probability indicates the most likely phase label for the user.
@@
TABLE I: Input data and outputs in topology problems
@@
[27] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Optimization based [16] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
@@
Heuristic Algorithms [28] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
@@
However, the key step in the above approaches is the stepwise subtraction of the identified customers or the users with the highest possibility labels, which ensures the accuracy of the phase labels of the customers who are located far away from the substation. Nevertheless, this key step might induce inaccuracies, since the accuracy of the current step is subject to that of the previous one. To address this issue, based on a modified genetic algorithm, a novel data-driven phase identification approach is introduced to identify the phase labels under incomplete datasets [28]. Based on multiple linear regression models, [27] presents a hybrid topology identification method to simultaneously identify line parameters and phase connection from raw SM data. Nevertheless, the underlying assumptions of the aforementioned methodologies—such as the presumption of complete dataSETS devoid of SM errors, and the omission of MV network impacts—significantly constrain their practical applicability [29]. Fig. 1 presents several underground feeder deployment patterns in the Netherlands [30], which indicates the challenges of LVDN topology identification and correction in practice, e.g., multiple cables under the same street. In addition, some of the above approaches are sensitive to missing data and measurement errors, e.g., SA-based approaches. Furthermore, the user-feeder connection identification issue is rarely considered in the aforementioned papers due to the assumption that one main feeder connects to the transformer.
To overcome these limitations, we propose a practical approach for topology correction in LVDNs using incomplete voltage data. Table I highlights the key differences between the proposed approach and existing approaches. By relying solely on voltage magnitude data, this approach preserves privacy and minimizes SM updates, as it releases the need for phase angle measurements from households. The corrected topology provides the foundational network visibility needed to implement load balancing, facilitate DERs integration, and alleviate LV line congestion, enabling active management. The main contributions are summarized as follows:
  • The operational states of the switch bars (including multiple switchgear) are encoded using Label Encoding rules, transforming the independent identification task of each switchgear into a classification problem that can be effectively addressed using existing supervised learning methods. In the correction analysis, a modified distance function is integrated into traditional Hierarchical Clustering (HC) to amplify the distinction between loads that are geographically close but belong to different phases or cables, thereby improving identification accuracy.
II Topology Correction approach
@@
Figure 3: Illustrative example of switch deployment and states in LVDNs: 1 represents the activated switch and vice versa [31].
II-A Switch State Identification
In LVDNs, multiple switch bars exist, each containing a different number of switches. An example of switch deployment is illustrated in Fig. 3. This non-uniform deployment poses challenges for accurately identifying the state of individual switches using SM data. To address this, a random forest (RF) algorithm is employed, leveraging voltage magnitude datasets to determine the state of switches within each switch bar. First, inspired by Label Encoding rules, the switch states in one bar are encoded into single-digital label formats, i.e., each combination of the switch states in the switch bar is uniquely mapped to a single digital label. The number of switches is assumed to be accessible for the DSO. For example, when there are 3 switches in the switch bar, the switch states $SS$ can be encoded as follows:
@@ II-A Switch State Identification
Through this selection process, the impact of the varying number of available SMs and class imbalance is minimized, leading to a more balanced and effective training dataset for the multi-class classification task. The selected voltage measurements can be represented in a matrix form $V_{train}$, forming the input for the multi-class classification task. Each row in matrix $V_{trian}$ represents training one sample with a encoded $SS$ label. $v^{s,C_{j}}_{m_{k},M}$ represents the voltage magnitude of the house $M$ connected to the feeder with $m_{k}$ houses when the switch state belongs to category $C_{j}$.
$\small V_{train}=\begin{bmatrix}v^{1,C_{1}}_{m_{1},1}&\cdots&v^{1,C_{1}}_{m_{1},M}&\cdots&v^{1,C_{1}}_{m_{k},1}&\cdots&v^{1,C_{1}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{S,C_{1}}_{m_{1},1}&\cdots&v^{S,C_{1}}_{m_{1},M}&\cdots&v^{S,C_{1}}_{m_{k},1}&\cdots&v^{S,C_{1}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{1,C_{J}}_{m_{1},1}&\cdots&v^{1,C_{J}}_{m_{1},M}&\cdots&v^{1,C_{J}}_{m_{k},1}&\cdots&v^{1,C_{J}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{S,C_{J}}_{m_{1},1}&\cdots&v^{S,C_{J}}_{m_{1},M}&\cdots&v^{S,C_{J}}_{m_{k},1}&\cdots&v^{S,C_{J}}_{m_{k},M}\\ \end{bmatrix}$ (4)
@@
The RF algorithm excels in multi-class classification by handling high-dimensional data, capturing complex feature interactions, and maintaining accuracy despite noise. Its ensemble learning approach—building diverse trees with random feature selection—prevents overfitting and enhances generalization. RF also provides feature importance insights, making it ideal for switch state identification. Given a dataset $D=\{(X_{1},y_{1}),\ldots,(X_{n},y_{n})\}$ where $X_{i}$ represents the feature vector (i.e., each row in $V_{trian}$) and $y_{i}$ is the encoded label by (1), the general process of RF training read as:
@@
    Draw $B$ bootstrap samples from the training data $V_{trian}$ and the corresponding labels.
@@
II-B User-feeder Connections Identification
The connection relations between customers and the feeders might be outdated and not always available for DSOs, which is ignored by multiple similar works [16, 19], especially for houses and multiple-layer buildings that are close to the streets with multiple deployed underground feeders. A practical case is shown in Fig. 4. Those unclear user-feeder connections also impact the phase label identifications. A correlation approach is introduced to reveal the user-feeder connections, which only rely on time-series voltage magnitude datasets. Two practical cases are addressed: 1) user-feeder connection identification without reference and 2) user-feeder connection identification with limited, accurate recordings.
II-B1 Connection Identification Without Recordings
The proposed user-feeder connection identification algorithm combines Hierarchical clustering (HC) algorithm and the correlation analysis based on the modified Pearson correlation coefficient (PCC). Algorithm 1 depicts the process of the proposed approach. The voltage magnitude dataset is stored in a matrix $V$, which is used as the input for the MPCC-based HC algorithm. Vector $V_{n}=[v_{n,1},v_{n,2},...,v_{n,\cal{T}}]^{T}$ represents the time series voltage magnitude at household $n$. $v_{n,t}$ represents the voltage magnitude at household $n$ at time $t$. $T$ is the measurement period of the SM data.
Input: $V$, $K_{s}$, $N$
1 for $n\leq N$ do
2       for $m\leq N$ do
3             $\text{PCC}(V_{n},V_{m})$ via Eq. (6)
4             $\text{FPCC}(\rho_{n,m})$ via Eq. (7)
5
6       end for
7
8 end for
9for $i\leftarrow 1$ to $N$ do
10       for $j\leftarrow i+1$ to $N$ do
11             $D_{ij}\leftarrow\text{MFP}(V_{n},V_{m})$ via Eq. (8)
12       end for
13
14 end for
15Obtain $\mathbf{D}\leftarrow\{D_{ij}\}$
16 Initialize $n$ clusters: $\mathcal{C}=\{C_{1},C_{2},...,C_{n}\}$
17 $t\leftarrow 0$
18 while $|\mathcal{C}|>1$ do
19       $(C_{a},C_{b})=\arg\min_{(C_{i},C_{j})\in\mathcal{C}}\max_{x_{i}\in C_{i},x_{j}\in C_{j}}D_{ij}$
20       $C_{\text{new}}\leftarrow\text{Merge }C_{a}\text{ and }C_{b}$
21       $D_{C_{\text{new}},C_{k}}\leftarrow\max(D_{C_{a},C_{k}},D_{C_{b},C_{k}})\quad\forall C_{k}\in\mathcal{C}$
22       $\mathbf{linkage\_matrix}[t]\leftarrow(C_{a},C_{b},D_{C_{a},C_{b}},\text{size}(C_{\text{new}}))$
23
24      Remove $C_{a},C_{b}$ from $\mathcal{C}$ and add $C_{\text{new}}$
25       $t\leftarrow t+1$
26 end while
27
28while $|\mathcal{C}|>K_{s}$ do
29       $(i,j)\leftarrow\text{find closest clusters via }\mathbf{linkage\_matrix}$
30       $C_{i}\leftarrow C_{i}\cup C_{j}$
31       Update $\mathcal{C}$
32
33 end while
34labels $\leftarrow$ Assigned labels to households based on $\mathcal{C}$
@@ 31       Update $\mathcal{C}$
$PCC(V_{n},V_{m})=\frac{\sum_{t=1}^{T}(v_{n,t}-\bar{V_{n}})(v_{m,t}-\bar{V_{m}})}{\sqrt{\sum_{t=1}^{T}(v_{n,t}-\bar{V_{n}})^{2}}\sqrt{\sum_{t=1}^{T}(v_{m,t}-\bar{V_{m}})^{2}}}$ (6)
@@
where $\alpha$ added to the denominator is used to avoid an infinite value of $FPCC$ and to control its distribution region.
To amplify the difference between voltage from different feeders, a modified $FPCC$ value (MFP) is employed to replace the Euclidean distance in the traditional HC algorithm.
@@
The second term in (8) is a modified likelihood function $F(\cdot)$ [26], which is used to calculate the $MFP(\cdot)$ between each sample $V_{n}$ and the rest of the samples. Since the range of the function $F(\cdot)$ is $(0,1]$, the range of the distance metric $MFP(\cdot)$ is correspondingly $[0,1)$. A value of $MFP(V_{n},V_{m})$ closer to 00 indicates a higher likelihood that the two samples $V_{n}$ and $V_{m}$ are collected from the same substation.
Given the switch state obtained from Section II-A, the feeder-transformer connection is available, and the number of feeders connected to the same transformer is assumed to be accessible for the DSO and taken as the number of clusters (i.e., $K_{s}$). Lines 1-13 in Algorithm 1 is to obtain the MFP of the input voltage magnitude dataset. Lines 16-23 are to find the link matrix of matrix $D$. From line 23 onwards, input data are clustered into a $K_{s}$ cluster using the obtained linkage matrix. If two users share the same label, they are connected to the same underground feeder, and vice versa.
@@
1 for  $n\geq N_{un}$ do
2       for $m\leq N_{k}$ do
3             $D_{0}(n,m)$leftarrow $D$($V_{n}$,$V_{m}$) via Eq. (8)
4       end for
5      $\hat{y}=\arg\max_{c}\sum_{n\in\mathcal{N}_{k}}\mathbb{I}(y_{n}=c)$
6       $L_{un}[n]$leftarrow $\hat{y}$
7
8 end for
@@ 6       $L_{un}[n]$leftarrow $\hat{y}$
II-B2 Connection Identification With Recordings
@@
  • Identified connections with high confidence from Section II-B1.
@@
Once the distances between the new sample and all labelled samples are computed, the $k$ nearest neighbors are selected. The final classification is determined using majority voting.
@@
II-C Phase Label Identification
@@
After the second step in the proposed framework (i.e., user-feeder connection identification in Section II-B), the third step is to identify the phase labels of the customer connection for each feeder, which is a typical 3-cluster classification problem. The phase identification approach is introduced by modifying Algorithm 1, i.e., removing Fisher $z$-transformation and using normalized voltage data $V^{*}$ as inputs.

### v2

# Conceptualisation, Methodology, Software, Validation, Writing - original draft

@@ Abstract
To correct the outdated and incomplete topology of low-voltage distribution networks (LVDNs) solely based on voltage magnitudes, a data-driven approach is developed based on machine learning algorithms and correlation analysis. Meanwhile, to address the similarity among smart meter (SM) data induced by distributed photovoltaic (PV) systems, a time-based SM data selection strategy is combined with the proposed correlation analysis. Unlike offline approaches, the proposed approach uses up-to-date voltage magnitudes to help distribution system operators determine switch states via supervised learning and refine user–feeder connections and customer phase labels through a modified hierarchical clustering algorithm. The feasibility and robustness of the proposed approach are validated using modified real-world LVDNs and multiple incomplete SM datasets collected from customers in the Netherlands. The results demonstrate that the time-based SM data selection strategy effectively mitigates its impact on phase identification, and the corrected topology not only improves network observability but also supports network operators in load balancing and PV consumption.
keywords:
[…]
Introduction
@@
Balanced distribution network contributes to reducing voltage violations, line congestion, power loss, etc. Most existing phase balancing and reconfiguration problems are formulated as mixed-integer optimization problems that depend on network topologies [10098964]. However, these topologies are often unavailable or inaccurate for distribution system operators (DSOs) due to outdated or missing data, such as limited phase labels [vanin2024phase]. The increasing uncertainty of distributed energy resources (DERs), including household photovoltaic (PV) systems, heating pumps etc., impacts the topology reconfiguration frequency and challenges the identification of low-voltage distribution network (LVDN) topology [10026490, 10347462, 10475702]. Moreover, existing topology identification methods rely on complete smart meter (SM) datasets, the availability of which is often limited due to privacy concerns [9696306, costa2022identification]. Synthetic networks and benchmark models presented in [birchfield2016grid, 2020Non] are insufficient to represent the diversity of European LVDNs for practical use by DSOs (e.g., state estimation). Thus, practical topology correction methods for LVDNs are needed for active management and DERs penetration.
@@
The topology identification problem in distribution networks (DNs) could be briefly classified into three categories based on the main issues they are aiming to solve: (1) switch state identification, (2) connection lines and their length identification, and (3) phase identification. The switch state identification was modeled as a classification problem in [9758816], solved by a trained deep neural network and providing quick estimations under the selected features. To recover accurate connection lines between transformers and customers, a probabilistic graphical model-based method was developed in [8456535] by restricting the propagation of SM data error. A power band-based strategy is proposed for dataset selection, aiming to enhance the accuracy of phase identification and transformer-to-meter matching [10533846]. In [liu2023hybrid], a hybrid data-driven approach integrating a partial correlation analysis strategy and a linear regression model is proposed to estimate topology from limited SM data. A similar work in [wang2023data] established a data-driven model based on the impedance matrix to correct topology and identify the impedance of connection lines from SM data.
@@
Based on an alternating direction method of multipliers, a robust topology identification method is proposed to estimate the line connection and their parameters using $\mu$-PMUs and SM datasets [9729419]. Based on incomplete and asynchronous SM data, an optimization-based data-driven approach is constructed to recover the approximate topologies in LVDNs with open datasets [liu2025topology]. A two-step approach is introduced to identify the topology of LVDNs with multiple meter types, which provide different resolution datasets, mitigating the dependence on the SMs [10400804]. For networks with low observability, a distance matrix is introduced in [10612997] to correct the topology and is also effective in detecting latent nodes positioned between observed nodes along network paths (i.e., intermediate latent nodes rather than endpoints). A physics-informed neural network (PINN) is constructed to estimate the topology by representing the nodes without meters, whose parameters are adjusted by a reinforcement learning-based approach [10138375]. Although the above methods generate accurate topologies of DNs, the requirements of complete SM data and prior topology information (e.g., the topology candidates) make them infeasible in practice due to data privacy-related concerns. Moreover, optimizing the line resistance and reactance as independent variables may lead to unrealistic estimation, as the properties are inherently interdependent and influenced by physical environments, such as temperature, humidity.
@@
Related to phase identification, current approaches rely on voltage magnitude data and demand profiles [7098424, garcia2025data]. Since the correlation of time-series voltage magnitude data from the same phase is stronger than that from different phases, correlation analysis is normally used. Given voltage magnitude datasets, spectral clustering and $k$-means clustering approaches are adopted to identify customer phase label [8955915]. The mixed integer linear programming model in [9409966] is constructed to identify the phase labels of single-phase customers by minimizing the gap between transformers’ measurements and the estimated data (e.g., power and voltage magnitude) in each phase. An optimization model-based approach was introduced in [10098964] to estimate customer phase labels in LVDNs with limited available recordings. Saliency analysis (SA) was integrated to extract the underlying features aiming to enhance the accuracy of phase label identification [7604110] SA and statistical tests are combined in [jimenez2020phase] to guarantee the identification accuracy of the data pairs with weak correlations. Voltage magnitude fluctuations among customers connected to different phases may exhibit similarities due to variations induced by the medium-voltage (MV) network or similar load patterns. To mitigate this effect, a high-pass filter is applied during time-series data pre-processing in the phase identification approach [hosseini2020machine]. Beyond hard classification (i.e., assigning customers a binary label of 0 or 1), soft classification is also considered for phase identification. The method in [garcia2023phase] assigns three probability values to customers. The highest probability indicates the most likely phase label for the user.
@@
Acronyms
MV/LV Medium/Low voltage LVDN Low-voltage distribution network DNs Distribution networks SM Smart meter DSO Distribution system operator DERs Distribution energy resources SA Saliency analysis PV Photovoltaic EV Electical vehicles HC Hierarchical clustering algorithm RF Random forest algorithm KNN K-nearest neighbors algorithm GNB Gaussian Naive Bayes algorithm SVM Support vector machine algorithm ANN Artificial neural network PINN Physics-informed neural network
        PCC Pearson correlation coefficient         MFP Modified PCC based Fisher z-transformation         GIS Geographic Information System         CN1 Network with two complex-structured feeders         CN2 Network with one simple- and one complex-structured feeder         SN Network with two simple-structured feeders         SNL Network with two longer simple-structured feeders         SNB Network with two simple-structured feeders with more branches         CC_PV CN1 with PV at each house         SC_PV SN with PV at each house         $C$ Complex feeder from CN1         SC Simple feeder from SN         CC_3_NU Complex feeder with three-phase non-uniform loads         SC_3_NU Simple feeder with three-phase non-uniform loads
However, the key step in the above approaches is the stepwise subtraction of the identified customers or users with the highest possibility labels, which ensures the accuracy of the phase labels of the customers who are located far away from the substation. Nevertheless, this key step might induce inaccuracies, since the accuracy of the current step is subject to that of the previous one. To address this issue, based on a modified genetic algorithm, a novel data-driven phase identification approach is introduced to identify the phase labels under incomplete datasets [jimenez2021new]. Based on multiple linear regression models, [cunha2020automated] presents a hybrid topology identification method to simultaneously identify line parameters and phase connection from raw SM data. Nevertheless, the underlying assumptions of the aforementioned methodologies—such as the presumption of complete datasets devoid of SM errors, and the omission of MV network impacts—significantly constrain their practical applicability [therrien2021assessment]. Fig. 1 presents several underground feeder deployment patterns in the Netherlands [web1], which indicates the challenges of LVDN topology identification and correction in practice, e.g., multiple cables under the same street. In addition, some of the above approaches are sensitive to missing data and measurement errors, e.g., SA-based approaches. Furthermore, the user-feeder connection identification issue is rarely considered in the aforementioned papers due to the assumption that one main feeder connects to the transformer.
To overcome these limitations, we propose a data-driven approach for topology correction in LVDNs using incomplete voltage data. Table 1 highlights the key differences between the proposed approach and existing approaches. By relying solely on voltage magnitude data, this approach preserves privacy and minimizes SM updates, as it releases the need for phase angle measurements from households. The corrected topology provides the foundational network visibility needed to implement load balancing, facilitate DERs integration, and alleviate LV line congestion, enabling active management. The main contributions are summarized as follows:
  • The operational states of the switch bars (including multiple switchgear) are encoded using Label Encoding rules, transforming the independent identification task of each switchgear into a classification problem that can be effectively addressed using existing supervised learning methods. In correlation analysis, a modified distance function is integrated into traditional Hierarchical Clustering (HC) to amplify the distinction between loads that are geographically close but belong to different phases or cables, thereby improving identification accuracy.
The remainder of this paper is organised as follows: Section 2 details the framework of the proposed approach, including identification of switch states, user-feeder connections, phase labels, and voltage data pre-processing. Section 3 presents the case studies and results. Finally, Section 4 summarizes the conclusions of this work.
Table 1: Input data and outputs in topology problems
@@
[cunha2020automated] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Optimization based [liu2025topology] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
@@
Heuristic Algorithms [jimenez2021new] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
@@
Topology Correction Approach
@@
Switch State Identification
In LVDNs, multiple switch bars exist, each containing a different number of switches. Each switch connects two cables (e.g., A and B), which are linked to different transformers (e.g., AT and BT). Under normal conditions, the switch remains open. When transformer AT becomes overloaded, the switch is closed, shifting the load on cable A to transformer BT. Therefore, the switch state plays a crucial role in maintaining network efficiency and enabling effective load shifting. An example of switch deployment is illustrated in Fig. 3. This non-uniform deployment poses challenges for accurately identifying the state of individual switches using SM data. To address this, a random forest (RF) algorithm is employed, leveraging voltage magnitude datasets to determine the state of switches within each switch bar. First, inspired by Label Encoding rules, the switch states in one bar are encoded into single-digital label formats, i.e., each combination of the switch states in the switch bar is uniquely mapped to a single digital label. The number of switches is assumed to be accessible for the DSO. For example, when there are 3 switches in the switch bar, the switch states $SS$ can be encoded as follows:
@@
Figure 3: Illustrative example of switch deployment and states in LVDNs: 1 represents the activated switch and vice versa [abb2025].
@@
Through this selection process, the impact of the varying number of available SMs and class imbalance is minimized, leading to a more balanced and effective training dataset for the multi-class classification task. The selected voltage measurements can be represented in a matrix form $V_{train}$, forming the input for the multi-class classification task. Each row in matrix $V_{train}$ represents training one sample with an encoded $SS$ label. $v^{s,C_{j}}_{m_{k},M}$ represents the voltage magnitude of the house $M$ connected to the feeder with $m_{k}$ houses when the switch state belongs to category $C_{j}$.
$V_{train}=\begin{bmatrix}v^{1,C_{1}}_{m_{1},1}&\cdots&v^{1,C_{1}}_{m_{1},M}&\cdots&v^{1,C_{1}}_{m_{k},1}&\cdots&v^{1,C_{1}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{S,C_{1}}_{m_{1},1}&\cdots&v^{S,C_{1}}_{m_{1},M}&\cdots&v^{S,C_{1}}_{m_{k},1}&\cdots&v^{S,C_{1}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{1,C_{J}}_{m_{1},1}&\cdots&v^{1,C_{J}}_{m_{1},M}&\cdots&v^{1,C_{J}}_{m_{k},1}&\cdots&v^{1,C_{J}}_{m_{k},M}\\ \vdots&&\vdots&&\vdots&&\vdots\\ v^{S,C_{J}}_{m_{1},1}&\cdots&v^{S,C_{J}}_{m_{1},M}&\cdots&v^{S,C_{J}}_{m_{k},1}&\cdots&v^{S,C_{J}}_{m_{k},M}\\ \end{bmatrix}$ (4)
@@
The RF algorithm excels in multi-class classification by handling high-dimensional data, capturing complex feature interactions, and maintaining accuracy despite noise. Its ensemble learning approach—building diverse trees with random feature selection—prevents overfitting and enhances generalisation. RF also provides feature importance insights, making it ideal for switch state identification. Given a dataset $D=\{(X_{1},y_{1}),\ldots,(X_{n},y_{n})\}$ where $X_{i}$ represents the feature vector (i.e., each row in $V_{train}$) and $y_{i}$ is the encoded label by (1), the general process of RF training read as:
@@
    Draw $B$ bootstrap samples from the training data $V_{train}$ and the corresponding labels.
@@
User-feeder Connections Identification
The connection relations between customers and the feeders might be outdated and not always available for DSOs, which is ignored by multiple similar works [liu2025topology, 10138375], especially for houses and multiple-layer buildings that are close to the streets with multiple deployed underground feeders. A practical case is shown in Fig. 4. Those unclear user-feeder connections also impact the phase label identifications. A correlation approach is introduced to reveal the user-feeder connections, which only rely on time-series voltage magnitude datasets. Two practical cases are addressed: 1) user-feeder connection identification without reference and 2) user-feeder connection identification with limited, accurate recordings.
Connection Identification Without Recordings
The proposed user-feeder connection identification algorithm combines the Hierarchical clustering (HC) algorithm and the correlation analysis based on the modified Pearson correlation coefficient (PCC). Algorithm 1 depicts the process of the proposed approach. The voltage magnitude dataset is stored in a matrix $V$, which is used as the input for the algorithm. Vector $V_{n}=[v_{n,1},v_{n,2},...,v_{n,\cal{T}}]^{T}$ represents the time series voltage magnitude at household $n$. $v_{n,t}$ represents the voltage magnitude at household $n$ at time $t$. $T$ is the measurement period of the SM data.
Input: $V$, $K_{s}$, NN
for $n\leq N$ do
 for $m\leq N$ do
    $\text{PCC}(V_{n},V_{m})$ via Eq. (6)
    $\text{FPCC}(\rho_{n,m})$ via Eq. (7)
   end for
end for
for $i\leftarrow 1$ to NN do
 for $j\leftarrow i+1$ to NN do
    $D_{ij}\leftarrow\text{MFP}(V_{n},V_{m})$ via Eq. (8)
   end for
end for
Obtain $\mathbf{D}\leftarrow\{D_{ij}\}$
Clusters $\mathcal{C}$: $\leftarrow$ Traditional HC and $\mathbf{D}$
labels $\leftarrow$ Assigned labels to households based on $\mathcal{C}$
@@
$PCC(V_{n},V_{m})=\frac{\sum_{t=1}^{T}(v_{n,t}-\bar{V_{n}})(v_{m,t}-\bar{V_{m}})}{\sqrt{\sum_{t=1}^{T}(v_{n,t}-\bar{V_{n}})^{2}}\sqrt{\sum_{t=1}^{T}(v_{m,t}-\bar{V_{m}})^{2}}}\quad$ (6)
@@
where $\alpha$ added to the denominator is used to avoid an infinite value of $FPCC(\cdot)$ and to control its distribution region.
To amplify the difference between voltage from different feeders, a modified $FPCC(\cdot)$ value (MFP) is employed to replace the Euclidean distance in the traditional HC algorithm.
@@
The second term in (8) is a modified likelihood function $F(\cdot)$ [garcia2023phase], which is used to calculate the $MFP(\cdot)$ between each sample $V_{n}$ and the rest of the samples. Since the range of the function $F(\cdot)$ is $(0,1]$, the range of the distance metric $MFP(\cdot)$ is correspondingly $[0,1)$. A value of $MFP(V_{n},V_{m})$ closer to 0 indicates a higher likelihood that the two samples $V_{n}$ and $V_{m}$ are collected from the same substation.
Given the switch state obtained from Section 2.1, the feeder-transformer connection is available, and the number of feeders connected to the same transformer is assumed to be accessible for the DSO and taken as the number of clusters (i.e., $K_{s}$). Lines 1-13 in Algorithm 1 are to obtain the MFP of the input voltage magnitude dataset. Lines 16-23 are to find the link matrix of matrix $D$. From line 23 onwards, input data are clustered into a $K_{s}$ cluster using the obtained linkage matrix. If two users share the same label, they are connected to the same underground feeder, and vice versa.
@@
for $n\geq N_{un}$ do
 for $m\leq N_{k}$ do
    $_{0}(n,m) \leftarrow DD$($V_{n}$$V_{m}$) via Eq. (8)
   end for
 $\hat{y}=\arg\max_{c}\sum_{n\in\mathcal{N}_{k}}\mathbb{I}(y_{n}=c)$
 $L_{un}[n]$ $\leftarrow$ $\hat{y}$
end for
@@
Connection Identification With Recordings
@@
  • Identified connections with high confidence from Section  2.2.1.
@@
Once the distances between the new sample and all labelled samples are computed, the $k$ nearest neighbours are selected. The final classification is determined using majority voting.
@@
Phase Label Identification
@@
Figure 6: Distribution of voltage magnitudes in a LVDN:(a) winter datasets and (b) summer datasets.
After the second step in the proposed framework (i.e., user-feeder connection identification in Section 2.2), the third step is to identify the phase labels of the customer connection for each feeder, which is a typical 3-cluster classification problem. The phase identification approach is introduced by modifying Algorithm 1, i.e., removing Fisher $z$-transformation and using normalized voltage data $V^{*}$ as inputs.